OpenStudy (gabbyalicorn):

Please help! I have no idea! D: The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 20, b = 28, c = ? A. 34.41 B. 35.45 C. 34.72 D. 35.88

OpenStudy (anonymous):

\[a^2+b^2=c^2\\ 20^2+28^2=c^2\]then take the square root to find c

OpenStudy (gabbyalicorn):

Okay, so 20^2+28^2 is 1184 then i find the square root which is 4?

OpenStudy (gabbyalicorn):

I definetly did something wrong.

OpenStudy (anonymous):

for sure the square root of 1184 is not 4!!!

OpenStudy (anonymous):

the 1184 is right

OpenStudy (anonymous):

now find \[\sqrt{1184}\]

OpenStudy (anonymous):

i would use a calculator

OpenStudy (gabbyalicorn):

How do I find the square roots, again?

OpenStudy (anonymous):

calculator i use this http://www.wolframalpha.com/input/?i=sqrt%28+20^2%2B28^2%29

OpenStudy (gabbyalicorn):

It shows up as 4, or maybe im not seeing this right.

OpenStudy (gabbyalicorn):

4√74

OpenStudy (anonymous):

yeah, but look at the part that has the decimal

OpenStudy (gabbyalicorn):

34.4?

OpenStudy (anonymous):

yeah actually \[34.409\] wihch would round to \[34.41\]

OpenStudy (gabbyalicorn):

just to summarize everything how did we get 34.41

OpenStudy (gabbyalicorn):

Like, what are the basic steps...I need to write this down x'D

OpenStudy (anonymous):

squared the length of each side added those numbers took the square root of the result

OpenStudy (anonymous):

https://www.youtube.com/watch?v=AuVi2ib-0qQ

OpenStudy (anonymous):

\[\sqrt{20^2+28^2} \]

rebeccaxhawaii (rebeccaxhawaii):

that is the formula

OpenStudy (gabbyalicorn):

the √20^2+28^2 or the squared the length of each side added those numbers took the square root of the result

OpenStudy (anonymous):

yes, they are the same

OpenStudy (anonymous):

square, add, take the square root, gets encapsulated as \[\sqrt{a^2+b^2}\]

OpenStudy (gabbyalicorn):

oooooh

OpenStudy (gabbyalicorn):

thank you!